# Biases in gendered citation practices: an exploratory study and some reflections on the Matthew and Matilda effects

### v1

@@ III-F Step 5: Statistical citation analysis
To simplify our citation analysis as in the literature (e.g., [5, 6, 9, 11, 47]), we only focus on the first and last authors of each paper available in our dataset. The last author is usually the most senior author in some fields [6, 47]. To support the citation analysis, Fulvio et al. [9] rely on an index called the Gender Citation Balance Index. The latter allows comparing the amount of citation each gender category receives with respect to the existing distribution of gender categories in SE literature. Equation 1 formalizes that index. To perform our citation analysis, we apply that equation separately to each gender category. A positive index yields a gender category that is more cited than expected (i.e. positive bias) [9]. A negative index yields a gender category that is less cited than expected (i.e. negative bias) [9].
@@ III-F2 Calculating observed proportion
The observed proportion (see equation 3) represents the percentage of citations which are authored by a specific gender category. This is known as our observed proportion, because the distribution of gender as it appears among citations in our dataset.
@@ III-G Step 6: bootstrapping
Like Fulvio et al. [9], we bootstrapped the 95% confidence interval for each gender category using 1000 iterations of random sampling with replacement from the total number of papers obtained in Step 4 i.e. on 96,282 papers. Thus, for each iteration, we computed the Gender Citation Balance Index for each category, considering: 1) the 2.5th percentile as the lower bound of the confidence interval; and 2) the 97.5th percentile as the upper bound of the confidence interval.
We performed these statistical steps once for the entire dataset, then again for each set of papers (i.e. subset of the dataset) associated with a gender category. This allows analyzing the citing behaviour of the following gender categories: MM, WW, and WuW. When calculating the Gender Citation Balance Indices for the subsets, the expected proportion remains the same from the full dataset. Additionally, each subset is filtered from the full dataset and bootstrapped individually (i.e. not subdivided from the already bootstrapped dataset).
@@ IV-B RQ2: temporal evolution of citations trends in SE journals
Recall from above that we obtained 251,174 citations when completing Step 4 of our methodology. We used these citations to plot the temporal evolution of citations trends in the SE literature. Figure 3 illustrates the corresponding chart. Thus, that Figure shows the observed citation proportions for each gender category from 2009 to 2024. Interestingly, when analysing that chart, we can notice that diversity in the SE literature seems to have slightly increased over time, with female-led papers (i.e. papers in the WW category) being increasingly more cited as time passes. But that increase has been rather marginal throughout the years since the citation proportion of the WW category has only increased by 1% from 2009 to 2024, while, as Figure 2 shows, the proportion of female authors has nearly doubled during the same period. Besides, as Figure 3 shows, the most cited category is the MM category, while the least cited category remains the WW category. Thus, time has not sufficiently witnessed fairness in gendered citation practices.
@@ IV-C1 Men usually tend to cite men more than women
our results show men usually cite men more often than women. Figure 4(b) illustrates that pattern. Thus, it shows the citation behavior of the MM category toward all the analyzed gender categories. That Figure shows the MM category usually over-cites the MM category. This pattern aligns with homophily [19], and is in accordance with the outcomes of Tahamtan et al. [29]’s review that concludes that male authors usually prefer to cite male authors over female authors. This pattern is also in accordance with the Matthew effect [24, 43]. The latter conveys the idea that research conducted by men is perceived as the most central and critical in a field [24, 43]. This pattern is also in accordance with the Matilda effect [28, 52]. The latter conveys the idea that the achievements of women are less acknowledged or research findings made by women are usually misallocated to other male authors [43, 28]. Noteworthy, the dominance of men in the SE journals authorship (see Figure 2) significantly amplifies their impact on gendered citation practices. Hence, since male authors are the ones citing the most, they are the ones who drive gendered citation practices.
@@ V Discussion and implications
Our results are similar to the ones the literature (e.g., [5, 6, 9, 11, 14, 29, 36, 39]) obtained and/or discussed in other disciplines (e.g., neuroscience, physics, and communication). Thus, gender disparities are pervasive in various disciplines including computing (software engineering). Accordingly, our results indicate that some efforts still need to be done to achieve fairness in citation practices in the SE field. As stated in Section II, such efforts may notably consist in the inclusion of CDSs in manuscripts submitted for publication in SE journals and conferences. Including CDSs in manuscripts could foster fairness in citations by promoting equal opportunities for researchers, and balancing social justice scales [37]. In many fields, the literature (e.g., [4, 7, 10, 11, 14, 19, 25, 51]) advocates for the inclusion of CDSs in manuscripts. Still, promoting diversity in reference lists should not be detrimental to the compliance with established standards of citation ethics [25].
@@
As Larivière et al. [1] point out, nearly half the population of each country is impacted by gender disparities in science and the lack of opportunities it yields for women. Citation biases, as observed in our study, exacerbate such disparities. To tackle gender disparities, it is therefore crucial for each country to also revisit its policies to enhance women’s participation in the scientific workforce [1]. Hence, in line with Lariviere et al. [1]’s conclusions, to avoid the repetition of past order that led to gender disparities, there is a need to integrate into policies the various local contexts (e.g. social, cultural, economic and political contexts) in which scientific activities are performed.
@@ VI Threats to Validity
As other existing approaches (e.g., [5, 6, 9, 11, 13, 14, 19, 43]), we also considered gender as a binary attribute. Hence, we did not take into account all the possible gender-related identities. This limitation stems from the capabilities of the gender assignment tools we used to determine the gender of authors. More specifically, these tools deal with the gender as a binary attribute. This does not sufficiently reflect the gender diversity of authors as some of them may not have a binary identity. Still, finding non-binary information is a very difficult task for studies centered on fairness analysis in general [32, 33, 34], especially since such information can be difficult to infer. However, to better foster diversity in citation practices, it is crucial to address that limitation in future work.
As in literature (e.g., [5, 6, 9, 11, 47]), we focused on the first and last authors of each paper in our dataset. This simplifies the citation analysis but does not allow capturing all the authorship nuances. Future work should focus on finding more efficient ways to capture authorship in citation analysis.
Our citation analysis only focuses on the gender of authors. This hinders the generalization of our study outcomes to other attributes and therefore to other under-represented groups. Thus, there is a need to also explore other attributes such as race, prolificity, disability, class, seniority, and citizenship, as well as their intersectionality with gender [6, 11, 14]. This could yield a more robust citation analysis. This could help create a broader database of under-represented scholars [11] and help devise solutions to foster fairness in citation practices.
@@ VII Conclusion and future work
Our results indicate that it is crucial to encourage journal editors and conference chairs to recommend the inclusion of citation diversity statements in manuscripts submitted for publication in software engineering journals and conferences. Developing tools that will automatically generate and include citation diversity statements in such manuscripts is also key to promote fairness in citation practices. This is in accordance with Bruton et al. [37] who advocate for the development of tools and initiatives allowing to improve citation practices.
@@
Finally, future work will study the impact that citation practices have on researchers’ careers, including recruitment, promotion, leadership, awards, publication rates, and opportunities for collaboration.
Abbreviated Citation Diversity Statement
In various disciplines, literature (e.g., [5, 6, 9, 11, 12, 20, 43]) concluded there are citations biases toward authors belonging to some socio-cultural groups (e.g., women, Black and Hispanic). We acknowledge that bias and have made some efforts to mitigate it when creating our reference list.

### v3

@@ III-F Step 5: Statistical citation analysis
To simplify our citation analysis as in the literature (e.g., [5, 6, 9, 11, 14, 47]), we only focus on the first and last authors of each paper available in our dataset. The last author is usually the most senior author in some fields [6, 47]. To support the citation analysis, Fulvio et al. [9] rely on an index called the Gender Citation Balance Index. The latter allows comparing the amount of citation each gender category receives with respect to the existing distribution of gender categories in SE literature. Equation 1 formalizes that index. To perform our citation analysis, we apply that equation separately to each gender category. A positive index yields a gender category that is more cited than expected (i.e. positive bias) [9]. A negative index yields a gender category that is less cited than expected (i.e. negative bias) [9].
@@ III-F2 Calculating observed proportion
The observed proportion (see equation 3) represents the percentage of citations which are authored by a specific gender category. It is known as our observed proportion, because it reflects the distribution of gender as it appears among citations in our dataset.
@@ III-G Step 6: bootstrapping
Like Fulvio et al. [9], we bootstrapped the 95% confidence interval for each gender category using 1000 iterations of random sampling with replacement from the total number of papers obtained in Step 4 i.e. 96,282 papers. Thus, for each iteration, we computed the Gender Citation Balance Index for each category, considering: 1) the 2.5th percentile as the lower bound of the confidence interval; and 2) the 97.5th percentile as the upper bound of the confidence interval.
We performed these statistical steps once for the entire dataset, then again for each set of papers (i.e. subset of the dataset) associated with a gender category. This allows analyzing the citing behaviour of the following gender categories: MM, WW, and WuW. When calculating the Gender Citation Balance Indices for the subsets, the expected proportion remains the same from the full dataset. Additionally, each subset is filtered out from the full dataset and bootstrapped individually (i.e. not subdivided from the already bootstrapped dataset).
@@ IV-B RQ2: temporal evolution of citations trends in SE journals
Recall from above that we obtained 251,174 citations when completing Step 4 of our methodology. We used these citations to plot the temporal evolution of citations trends in the SE literature. Figure 3 illustrates the corresponding chart. Thus, that Figure shows the observed citation proportions for each gender category from 2009 to 2024. Interestingly, when analyzing that chart, we can notice that diversity in the SE literature seems to have slightly increased over time, with female-led papers (i.e. papers in the WW category) being increasingly more cited as time passes. But that increase has been rather marginal throughout the years since the citation proportion of the WW category has only increased by 1% from 2009 to 2024, while, as Figure 2 shows, the proportion of female authors has nearly doubled during the same period. Besides, as Figure 3 shows, the most cited category is the MM category, while the least cited category remains the WW category. Thus, time has not sufficiently witnessed fairness in gendered citation practices.
@@ IV-C1 Men usually tend to cite men more than women
our results show men usually cite men more often than women. Figure 4(b) illustrates that pattern. Thus, it shows the citation behavior of the MM category toward all the analyzed gender categories. That Figure shows the MM category usually over-cites the MM category. This pattern aligns with homophily [19], and is in accordance with the outcomes of Tahamtan et al. [29]’s review that concludes that male authors usually prefer to cite male authors over female authors. This pattern is also in accordance with the Matthew effect [24, 43]. The latter conveys the idea that research conducted by men is perceived as the most central and critical in a field [24, 43]. This pattern is also in accordance with the Matilda effect [28, 52]. The latter conveys the idea that the achievements of women are less acknowledged or research findings made by women are usually misallocated to other male authors [43, 28]. Note that, the dominance of men in the SE journals authorship (see Figure 2) significantly amplifies their impact on gendered citation practices. Hence, since male authors are the ones citing the most, they are the ones who drive gendered citation practices.
@@ V Discussion and implications
Our results are similar to the ones the literature (e.g., [5, 6, 9, 11, 14, 29, 36, 39]) obtained and/or discussed in other disciplines (e.g., neuroscience, physics, communication). Thus, gender disparities are pervasive in various disciplines including computing (SE) and are the fruit of longstanding systemic issues. Accordingly, our results indicate that some efforts need to be done to achieve fairness in citation practices in the SE field. As stated in Section II, such efforts may consist in the inclusion of CDSs in manuscripts submitted for publication to SE journals and conferences. Including CDSs in manuscripts could foster fairness in citations by promoting equal opportunities for researchers, and balancing social justice scales [37]. In many fields, the literature (e.g., [4, 7, 10, 11, 14, 19, 25, 51]) advocates for the inclusion of CDSs in manuscripts. Still, promoting diversity in reference lists should not be detrimental to the compliance with established standards of citation ethics [25].
@@
Nearly half the population of each country is impacted by gender disparities in science and the lack of opportunities it yields for women [1]. Citation biases, as observed in our study, exacerbate such disparities. To tackle gender disparities, it is therefore crucial for each country to also revisit its policies to enhance women’s participation in the scientific workforce [1]. Hence, in line with Lariviere et al. [1]’s conclusions, to avoid the repetition of past order that led to gender disparities, we contend there is a need to integrate into policies the various local contexts (e.g. social, cultural, economic and political contexts) in which scientific activities are performed.
@@ VI Threats to Validity
As other existing approaches (e.g., [5, 6, 9, 11, 13, 14, 19, 43]), we also considered gender as a binary attribute. Hence, we did not take into account all the possible gender-related identities. This limitation stems from the capabilities of the gender assignment tools we used to determine the gender of authors. More specifically, these tools deal with the gender as a binary attribute. This does not sufficiently reflect the gender diversity of authors as some of them may not have a binary identity. Still, finding non-binary information is a very difficult task for studies centered on fairness analysis in general [32, 33, 34, 35], especially since such information can be difficult to infer. However, to better foster diversity in citation practices, it is crucial to address that limitation in future work.
As in most approaches in the literature (e.g., [5, 6, 9, 11, 47]), and in the approaches we adapted (e.g., [9, 11]) as well as their supporting tools, our citation analysis focuses on the first and last authors of each paper in our dataset. This simplifies the citation analysis but does not allow capturing all the authorship nuances. Future work should focus on finding more efficient ways to capture authorship in citation analysis.
Our analysis only focuses on the gender of authors. This hinders the generalization of our study outcomes to other attributes and/or under-represented groups. Thus, there is a need to also explore other attributes such as race, prolificity, disability, class, seniority, and citizenship, as well as their intersectionality with gender [6, 11, 14]. This could yield a more robust citation analysis. This could also help create a broader database of under-represented scholars [11] and help devise solutions to foster fairness in citation practices.
@@ VII Conclusion and future work
Our results suggest that it is crucial to encourage journal editors and conference chairs to recommend the inclusion of citation diversity statements in manuscripts submitted for publication in software engineering journals and conferences. Developing tools that will automatically generate and include citation diversity statements in such manuscripts is also key to promote fairness in citation practices. This is in accordance with Bruton et al. [37] who advocate for the development of tools and initiatives allowing to improve citation practices.
@@
Future work will also study the impact of citation practices on researchers’ careers (eg., recruitment, promotion, leadership, awards, publication rates, collaboration opportunities).
Full Citation Diversity Statement
In various disciplines, literature (e.g., [5, 6, 9, 11, 12, 20, 43]) concluded there are citation biases toward authors belonging to some socio-cultural groups (e.g., women, Black and Hispanic). We acknowledge these biases and, in accordance with citation ethics [37], we have made some efforts to mitigate them when creating our reference list. With respect to gender categories, our reference list (excluding self-citations) contains 27.45% man/man, 17.65% man/woman, 29.41% woman/man, 11.76% woman/woman and 13.73% unknown gender authorship. With respect to race categories, our reference list (excluding self-citations) contains 64.71% white/white, 5.88% white/author of colour, 17.65% author of colour/white, 5.88% author of colour/author of colour and 5.88% unknown race authorship. We performed gender and race classifications using respectively Gender-API and Ethnicolr’s Florida Voter Registration Model. These tools either deal with gender as a binary attribute or consider limited racial identities. Future tools should address such limitations.
Acknowledgment
We thank the LURA (Lassonde Undergraduate Research Award) program for funding this research.